Calculate the magnitudes of magnetic intensity H and the magnetic field
B at the centre of a 1500-turn solenoid which is 0.22 m long and carries a current of 1.5 A.
(µ0 =4π×10^-7H/m)
1
Expert's answer
2020-03-09T10:53:22-0400

Magnetic induction at the center of the solinoid is

"B=\frac{\mu_0 \cdot I \cdot N}{l}"

We have

"N=1500"

"I=1.5A" "l=0.22 m"

Then

"B=\frac{\mu_0 \cdot I \cdot N}{l}=\frac{4\pi \cdot 10^{-7}\cdot 1.5 \cdot 1500}{0.22}=4\pi \cdot 10^{-7}\cdot1.023 \cdot10^{4}=12.85mT"

Magnetic intensity and magnetic induction are related by the formula

"H=\frac{B}{\mu_0}"

The magnetic intensity is

"H=\frac{B}{\mu_0}=\frac{4\pi \cdot 10^{-7}\cdot1.023 \cdot10^{4}}{4\pi \cdot 10^{-7}}=1.023 \cdot10^{4} A/m"